# BESSTIE: A Benchmark for Sentiment and Sarcasm Classification for Varieties of English

### v1

@@
Abstract.
Despite large language models (LLMs) being known to exhibit bias against non-mainstream varieties, there are no known labeled datasets for sentiment analysis of English. To address this gap, we introduce BESSTIE, a benchmark for sentiment and sarcasm classification for three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK). Using web-based content from two domains, namely, Google Place reviews and Reddit comments, we collect datasets for these language varieties using two methods: location-based and topic-based filtering. Native speakers of the language varieties manually annotate the datasets with sentiment and sarcasm labels. Subsequently, we fine-tune nine large language models (LLMs) (representing a range of encoder/decoder and mono/multilingual models) on these datasets, and evaluate their performance on the two tasks. Our results reveal that the models consistently perform better on inner-circle varieties (i.e., en-AU and en-UK), with significant performance drops for en-IN, particularly in sarcasm detection. We also report challenges in cross-variety generalisation, highlighting the need for language variety-specific datasets such as ours. BESSTIE promises to be a useful evaluative benchmark for future research in equitable LLMs, specifically in terms of language varieties. The BESSTIE datasets, code, and models are currently available on request, while the paper is under review. Please email aditya.joshi@unsw.edu.au.
bias, language variety, dialectal bias, language models, sentiment classification, sarcasm classification, benchmark, natural language processing
ccs: Computing methodologies Natural language processingccs: Computing methodologies Language resources
1. Introduction
Benchmark-based evaluation (Socher et al., 2013) of large language models (LLMs) is the prevalent norm in natural language processing (NLP). Benchmarks provide labeled datasets (Wang et al., 2019; Muennighoff et al., 2023) for specific tasks so that LLMs can be evaluated against them. However, most NLP benchmarks do not contain text in language varieties such as national varieties, dialects, sociolects or creoles (Plank, 2022; Lent et al., 2024; Joshi et al., 2024). As a result, despite the superlative performance of LLMs on Standard American English, LLMs are not evaluated on other varieties of English. Knowing that LLMs can be biased against certain language varieties (as shown by  (Deas et al., 2023) for African-American English), little empirical evidence is found for their bias towards several language varieties, particularly those of English. An exception is Multi-VALUE (Ziems et al., 2023) which creates synthetic datasets of language varieties using linguistically-informed syntactic transformations and reports a degraded performance on the GLUE benchmark. However, such synthetic texts are insufficient in measuring language variety bias because language varieties are more than syntax and encompass orthography, vocabulary, and cultural pragmatics (Nguyen, 2021). As a result, labeled datasets of natural text are crucial to measure the bias in LLMs towards these varieties.
Benchmark Sent. Sarc. Eng. Var.
Cieliebak et al. (2017) (Cieliebak et al., 2017)
Wang et al. (2018) (Wang et al., 2018)
Alharbi et al. (2020) (Alharbi et al., 2020)
Abu Farha et al. (2021) (Abu Farha et al., 2021)
Elmadany et al. (2023) (Elmadany et al., 2023)
Faisal et al. (2024) (Faisal et al., 2024)
@@ 1. Introduction
Table 1. Comparison of BESSTIE with past benchmarks for sentiment or sarcasm classification. ‘Sent.’ indicates sentiment classification, ‘Sarc.’ denotes sarcasm classification, ‘Eng.’ denotes English, and ‘Var.’ denotes language varieties. A checkmark (✓) denotes the availability of a particular feature, while a cross (✕) indicates its absence.
To address this gap, we introduce BESSTIE, A BEnchmark for Sentiment and Sarcasm classification for varieTIes of English, namely, Australian (en-AU), Indian (en-IN), and British (en-UK). BESSTIE comprises a manually labeled dataset of textual posts along with performance results of fine-tuned LLMs on the two classification tasks. The dataset consists of text samples collected from two web-based domains, namely, Google Places reviews, and Reddit comments, using two filtering methods: location-based and topic-based filtering respectively. Following that, we obtain manual annotations for boolean sentiment and sarcasm labels from native speakers of the language varieties. The novelty of the BESSTIE dataset can be seen in Table 1. BESSTIE improves not only upon popular sentiment/sarcasm-labeled datasets but also a recent dialectal benchmark, DialectBench (Faisal et al., 2024), which contains dialectal datasets for sentiment classification but for languages other than English. Therefore, the BESSTIE dataset is novel in the introduction of new language varieties (specifically, varieties of English) and, an additional task, sarcasm classification.
We then use the dataset to model both the tasks (sentiment and sarcasm classification) as binary classification problems. We evaluate nine fine-tuned LLMs, spanning encoder/decoder and monolingual/multilingual models. Our results indicate that current models exhibit a significant decrease in performance when processing text written in en-IN, an outer-circle variety, in comparison with inner-circle varieties of English [^1] Inner-circle English varieties are those spoken by people who use English as a first language. Outer-circle varieties are those spoken by bi/multilingual speakers. (en-AU and en-UK). The performance on sarcasm classification is consistently low across the three varieties, providing pointers for future work. Following that, we examine the aggregate performance for the three varieties, the two tasks and model attributes (encoder/decoder and monolingual/multilingual). We also examine cross-variety and cross-domain performance on all models highlighting the need for BESSTIE. BESSTIE evaluation provides a solid baseline to examine biases of contemporary LLMs, and will help to develop new methods to mitigate these biases.
The contributions of this work are:
  • Creation of a manually annotated dataset of user-generated content with sentiment and sarcasm labels for three varieties of English, called BESSTIE.
  • Evaluating nine LLMs on BESSTIE and analysing their performance to identify challenges arising due to language varieties.
An overview of BESSTIE is described in Figure 1. The paper describes the dataset (Section 4; referring to ‘Collection’ and ‘Annotation’ in the figure) and the evaluation methodology (Section 5; ‘Evaluation’ in the Figure) of BESSTIE. The experiment setup and results are in Sections 6 and 7 respectively, followed by error analysis in Section 8.
2. Related work
NLP Benchmarks such as GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and DynaBench (Kiela et al., 2021) have played a pivotal role in evaluating LLMs on multiple NLP tasks, including sentiment classification. However, these benchmarks are limited in their ability to capture the linguistic nuances of non-standard language varieties. Efforts to address this include datasets for African languages and dialects (Muhammad et al., 2023), Arabic dialects (Elmadany et al., 2023), and CreoleVal (Lent et al., 2023) which provides datasets for 28 creole languages. More recently, DIALECTBENCH (Faisal et al., 2024) is a collection of dialectal datasets, and associated NLP models across 281 dialectal varieties for several tasks. However, the benchmark does not contain English dialectal datasets for sentiment and sarcasm classification. BESSTIE is the first benchmark for sentiment and sarcasm classification specifically for non-standard varieties of English.
Similarly, several sarcasm classification datasets for standard variety of English exist, consisting of reddit comments (Khodak et al., 2018), amazon reviews (Filatova, 2012), and tweets (Ptáček et al., 2014; Painter et al., 2022; Abercrombie and Hovy, 2016; Oprea and Magdy, 2020). However, there is a notable absence of sarcasm classification datasets that account for non-standard varieties of English. ArSarcasm-v2 (Abu Farha et al., 2021) is a dataset annotated for sarcasm, sentiment, and the dialect of each tweet for Arabic. Abu Farha and Magdy (2021) benchmarked Arabic-specific transformer models on this dataset, noting that models trained on Arabic data, including varieties of the language, tend to perform better.
We emphasise that the papers listed above approach benchmark creation in a similar manner as BESSTIE: they create an annotated dataset for a task, and test a broad range of models to highlight the potential for future work. In that sense, these benchmark papers acknowledge that creating a dataset and measuring bias is an important task achieved by papers that introduce benchmarks, without necessarily presenting a new method itself for the task. This holds true for our paper as well.
3. Ethical Considerations
The project received ethics approval from the Human Research Ethics Committee at UNSW, Sydney (reference number: iRECS6514). We treat the reviews as separate texts and do not aggregate them by user in any way. The Reddit comments and Google reviews used in this paper are from publicly available posts, accessible through official APIs. We adhere to the terms of service and policies outlined by these platforms while gathering data. We do not attempt to de-identify any users or collect aggregate information about individual users. The dataset, including sentiment and sarcasm annotations, is shared in a manner that preserves user privacy and abides by the rules and guidelines of the source platforms.
@@ 3. Ethical Considerations
Figure 1. Methodology of BESSTIE. (a) Collection is detailed in Section 4.1, (b) Annotation is detailed in 4.2, and (c) Evaluation is detailed in Section 5.
4. Dataset Creation
@@
We now describe parts (a) and (b) of Figure 1 which result in a manually annotated dataset for sentiment and sarcasm classification.
4.1. Data Collection
@@
We collect textual posts from two domains: Google Places reviews and Reddit comments, using two filtering methods: location-based and topic-based respectively. Location-based filtering implies that we select reviews that were posted for locations in the three countries. Specifically, we collect reviews posted in cities in AU, IN and UK, and their corresponding ratings ($1$ to $5$ stars, where $5$ is the highest.) using the Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview; Accessed on 23 July 2024 from the place types, also defined by the API. The criteria for city selection are based on population thresholds specific to each country (en-AU: $20$K; en-IN:$100$K; en-UK: $50$K). We further filter the dataset to only contain reviews with closer ratings ($2$ and $4$ stars). We also filter out any non-English content using language probabilities calculated by fastText (Grave et al., 2018) word vectors, using a threshold of $0.98$. Additionally, to mitigate the risk of including reviews written by tourists rather than residents, we exclude reviews from locations designated as ‘tourist attractions’ [^3] https://developers.google.com/maps/documentation/places/web-service/supported_types; Accessed on 13 November 2024. by the Google Places API.
@@
To create the reddit subset, we employ topic-based filtering and choose up to four subreddits per variety (en-AU: ‘melbourne’, ‘AustralianPolitics’,‘AskAnAustralian’; en-IN: ‘India’, ‘IndiaSpeaks’, ‘BollyBlindsNGossip’; en-UK: ‘England’, ‘Britain’, ‘UnitedKingdom’, ‘GreatBritishMemes’). We select these subreddits based on the understanding that they feature popular, discussions specific to a variety, making it highly likely that contributors use the appropriate language variety. For each variety, we scrape $12,000$ comments evenly across the selected subreddits, capping at $20$ comments per post and focusing on recent posts. These are then randomly sampled and standardized to $3,000$ comments per variety before manual annotation.
4.2. Annotation
We hire three annotators [^4] Two of these are the authors of the paper who were also involved in the evaluation. each for the three language varieties. The annotators assign the processed reviews and posts two labels: sentiment and sarcasm. The choice of labels given are negative, positive, and discard. We instruct the annotators to use the discard label for uninformative examples with no apparent polarity and in a few cases, for computer-generated messages. We compensate the annotators at the casual employment rate prescribed by the host institution.
@@ 4.2. Annotation
Figure 2. Confusion matrix showing the overlap in variety annotations between annotators, ant-IN, and ant-UK. Rows represent the labels assigned by ant-IN, and columns represent the labels assigned by ant-UK. The principle diagonal elements indicate agreement between annotators, while off-diagonal elements highlight disagreements.
4.3. Data Quality & Statistics
Data Quality
@@ 4.3. Data Quality & Statistics
While previous studies utilise location-based filtering to evaluate the robustness of language technologies across varieties (Blodgett et al., 2016; Sun et al., 2023), it needs to be examined if the collected text is indeed from these national varieties. To address this, we conduct an annotation exercise with en-IN and en-UK annotators, asking them to manually identify the variety of a given text. Specifically, they label a random sample of $300$ texts ($150$ from reviews and $150$ from comments) as en-AU, en-IN, en-UK, or Cannot say. Annotator agreement is measured using Cohen’s kappa ($\kappa$).

### v3

@@
Abstract
Despite large language models (LLMs) being known to exhibit bias against non-standard language varieties, there are no known labelled datasets for sentiment analysis of English. To address this gap, we introduce besstie, a benchmark for sentiment and sarcasm classification for three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK). We collect datasets for these language varieties using two methods: location-based for Google Places reviews, and topic-based filtering for Reddit comments. To assess whether the dataset accurately represents these varieties, we conduct two validation steps: (a) manual annotation of language varieties and (b) automatic language variety prediction. Native speakers of the language varieties manually annotate the datasets with sentiment and sarcasm labels. We perform an additional annotation exercise to validate the reliance of the annotated labels. Subsequently, we fine-tune nine large language models (LLMs) (representing a range of encoder/decoder and mono/multilingual models) on these datasets, and evaluate their performance on the two tasks. Our results show that the models consistently perform better on inner-circle varieties (i.e., en-AU and en-UK), in comparison with en-IN, particularly for sarcasm classification. We also report challenges in cross-variety generalisation, highlighting the need for language variety-specific datasets such as ours. besstie promises to be a useful evaluative benchmark for future research in equitable LLMs, specifically in terms of language varieties. The besstie dataset is publicly available at: https://huggingface.co/datasets/unswnlporg/BESSTIE.
BESSTIE: A Benchmark for Sentiment and Sarcasm Classification for Varieties of English
Dipankar Srirag1 Aditya Joshi1 Jordan Painter2 Diptesh Kanojia2 1University of New South Wales, Sydney, Australia 2Institute for People-Centered AI, University of Surrey, Surrey, United Kingdom {d.srirag,aditya.joshi}@unsw.edu.au{jp1106,d.kanojia}@surrey.ac.uk
1 Introduction
Benchmark-based evaluation Socher et al. (2013) of large language models (LLMs) is the prevalent norm in natural language processing (NLP). Benchmarks provide labelled datasets Wang et al. (2019); Muennighoff et al. (2023) for specific tasks so that LLMs can be evaluated against them. However, most NLP benchmarks do not contain text in language varieties such as national varieties, dialects, sociolects or creoles Plank (2022); Lent et al. (2024); Joshi et al. (2025).
@@ 1 Introduction
Table 1: Comparison of besstie with past benchmarks for sentiment or sarcasm classification. ‘Sent.’ indicates sentiment classification, ‘Sarc.’ denotes sarcasm classification, ‘Eng.’ denotes English, and ‘Var.’ denotes language varieties. A checkmark (✓) denotes the availability of a particular feature, while a cross (✕) indicates its absence.
@@
Figure 1: Creating dataset for besstie; Collection in Section 2.1, Quality Assessment in Section 2.2, Annotation in 2.3.
As a result, despite the superlative performance of LLMs on Standard American English, LLMs are not evaluated on other varieties of English. Knowing that LLMs can be biased against certain varieties of English, as shown by Deas et al. (2023) and Srirag et al. (2025) for African-American English and Indian English respectively, little empirical evidence is found for their bias towards several other varieties. An exception is Multi-VALUE Ziems et al. (2023) which creates synthetic datasets of language varieties using linguistically-informed syntactic transformations and reports a degraded performance on the GLUE benchmark. However, such synthetic texts do not accurately represent the spoken or written forms of the language variety because language varieties are more than syntax and encompass orthography, vocabulary, and cultural pragmatics Nguyen (2021). As a result, labelled datasets of natural text are crucial to measure the bias in LLMs towards these varieties.
Acknowledging the role of the sentiment classification task Socher et al. (2013) in GLUE, we introduce besstie, A BEnchmark for Sentiment and Sarcasm classification for varieTIes of English, namely, Australian (en-AU), Indian (en-IN), and British (en-UK). The relationship between sentiment and sarcasm classification as related tasks is well-understood in the context of sentiment analysis Chauhan et al. (2020). besstie comprises a manually labelled dataset of textual posts along with performance results of fine-tuned LLMs on the two classification tasks. The dataset consists of text samples collected from two web-based domains, namely, Google Places reviews, and Reddit comments, using two filtering methods: location-based and topic-based filtering respectively. We validate the representation of these varieties through a two-step process: (a) a manual annotation exercise and (b) automatic variety identification using models fine-tuned on the ICE-Corpora Greenbaum and Nelson (1996). Following that, we obtain manual annotations for boolean sentiment and sarcasm labels from native speakers of the language varieties.
We then use the dataset to model both the tasks as binary classification problems. We evaluate nine fine-tuned LLMs, spanning encoder/decoder and monolingual/multilingual models. Our results indicate that current models exhibit a significant decrease in performance when processing text written in en-IN, an outer-circle variety, in comparison with inner-circle varieties of English [^1] Inner-circle English varieties are those spoken by people who use English as a first language. Outer-circle varieties are those spoken by bi/multilingual speakers. (en-AU and en-UK). The performance on sarcasm classification is consistently low across the three varieties, providing pointers for future work. Following that, we examine the aggregate performance for the three varieties, the two tasks, and model attributes (encoder/decoder and monolingual/multilingual). We also examine cross-variety performance on all models highlighting the need for besstie. besstie evaluation provides a solid baseline to examine biases of contemporary LLMs, and will help to develop new methods to mitigate these biases. The novelty of the besstie dataset can be seen in Table 1. besstie improves not only upon popular sentiment/sarcasm-labelled datasets but also a recent dialectal benchmark, DialectBench Faisal et al. (2024), which covers sentiment classification but for dialects of languages other than English. Therefore, the besstie dataset is novel in the introduction of new language varieties (specifically, varieties of English) for both sentiment and sarcasm classification.
The contributions of this work are: (a) Creation of a manually annotated dataset of user-generated content with sentiment and sarcasm labels for three varieties of English, called besstie; (b) Evaluation of nine LLMs on besstie and analysing their performance to identify challenges arising due to language varieties.
2 Dataset Creation
@@
Figure 1 shows the process followed to create a manually annotated dataset for sentiment and sarcasm classification.
2.1 Data Collection
@@
We collect textual posts from two domains: Google Places reviews and Reddit comments, using two filtering methods: location-based and topic-based respectively. Location-based filtering implies that we select reviews that were posted for locations in the three countries. Specifically, we collect reviews posted in cities in AU, IN and UK, and their corresponding ratings ($1$ to $5$ stars, where $5$ is the highest.) using the Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview; Accessed on 10 February 2025 from the place types, also defined by the API. The criteria for city selection are based on population thresholds specific to each country (en-AU: $20$K; en-IN:$100$K; en-UK: $50$K). We filter out any non-English content using language probabilities calculated by fastText Grave et al. (2018) word vectors, using a threshold of $0.98$. Additionally, to mitigate the risk of including reviews written by tourists rather than residents, we exclude reviews from locations designated as ‘tourist attractions’ by the Google Places API.
[…]
Table 2: Performance of distil on sentiment classification with different labelling schemes. naive refers to labels based on 1 and 5 stars, while ours corresponds to labels based on 2 and 4 stars. $\mu$ denotes the average F-Score across all varieties
We also conduct a preliminary investigation using DistilBERT-Base (distil;Sanh et al., 2020) to examine the influence of label semantics, the star ratings, on the performance of sentiment classification. We experiment with distil on the said task using the collected Google Places reviews under two distinct labelling schemes: (a) extreme ratings, i.e., 1 and 5 stars (naive); and (b) moderate ratings, i.e., 2 and 4 stars (ours). Table 2 presents the performance of distil for each variety under both schemes, reported using the macro-averaged F-Score. We observe that distil yields strong performance when trained on naive labels, achieving an average $\mu$ of 0.96 across the three varieties. However, task performance drops considerably with the use of the more ambiguous ours labels, highlighting that they introduce increased difficulty and nuanced content. Following these findings, we filter the dataset to retain only reviews with intermediate ratings (2 and 4 stars), to better capture this nuance and avoid model overfitting to well-distinguished and polarised examples.
@@
To create the reddit subset, we employ topic-based filtering and choose up to four subreddits per variety (en-AU: ‘melbourne’, ‘AustralianPolitics’,‘AskAnAustralian’; en-IN: ‘India’, ‘IndiaSpeaks’, ‘BollyBlindsNGossip’; en-UK: ‘England’, ‘Britain’, ‘UnitedKingdom’, ‘GreatBritishMemes’), where the topics are determined by native speakers of these language varieties. We select these subreddits based on the understanding that they feature discussions specific to a variety, making it highly likely that the post authors use the corresponding language variety. For each variety, we scrape $12,000$ comments evenly across the selected subreddits, capping at $20$ comments per post and focusing on recent posts. These are then randomly sampled and standardised to $3,000$ comments per variety before manual annotation. We discard information such as user identifiers and post identifiers to maintain the anonymity of the user.
@@
Figure 2: Confusion matrix showing the overlap in variety annotations between annotators, ant-IN, and ant-UK. Rows represent the labels assigned by ant-IN, and columns represent the labels assigned by ant-UK. The principle diagonal elements indicate agreement between annotators, while off-diagonal elements highlight disagreements.
2.2 Quality Assessment
While previous studies utilise location-based filtering Blodgett et al. (2016); Sun et al. (2023), we examine if the collected text from the two methods is indeed from these language varieties. To address this, we conduct two evaluations: (a) Manual annotation for national variety identification, and (b) Automated validation of national varieties.
Manual annotation
@@ 2.2 Quality Assessment
We request en-IN and en-UK annotators, asking them to manually identify the variety of a given text. Specifically, they label a random sample of $300$ texts ($150$ from reviews and $150$ from comments) as en-AU, en-IN, en-UK, or Cannot say. Annotator agreement is measured using Cohen’s kappa ($\kappa$).